# Universal quantum control over non-Hermitian continuous-variable systems

### v1

@@ IV.2 Unidirectional perfect absorber
In this section, we show that the cavity-magnonic system under our UQC theory in Sec. IV.1 can simulate a nonreciprocal behavior under the same control conditions as in Fig. 3 for $\mu_{1}^{\dagger}(t)$ in the ket space. For example, when the Fock state is initially prepared in the magnon mode, i.e., $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$, the total population of the system is found to be nearly vanishing at the end of time evolution. Note the description about such a phenomenon is meaningful by our theory with no artificial renormalization. Moreover, this unidirectional absorbtion can be interpreted by the coupled-mode theory Xia et al. (2025) about coherent perfect absorber (CPA) Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025) to a certain degree.
@@
$=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}\begin{pmatrix}-i\gamma_{a}&i[J(t)+\Gamma]\\ -i[J(t)-\Gamma]&-i\gamma_{a}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$.
@@
where $S_{21}(\tau)\approx 4\gamma_{1}J(\tau)/\gamma_{a}^{2}\neq 0$. Here the second equivalence holds when $\gamma_{1}=\gamma_{a}/2$. Equation (41) indicates that for a nonvacuum initial state in the magnon mode $b$, the total system population can become almost zero at a desired moment, i.e., $a(\tau),b(\tau)]^{T}=S(t=\tau)[0,b(0)]^{T}\approx(0,0)^{T}$. In practice, the non-Hermitian cavity magnonic system now becomes a promising candidate Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) for the unidirectional perfect absorber under our UQC framework.
Figure 5: Fidelity dynamics about the individual states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ with various coefficients $\lambda$, under a $\mathcal{PT}$-symmetric-broken Hamiltonian. Inset: Numerical results for the logarithm of the nonreciprocity $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]$ as a function of $\lambda$. The other parameters are the same as Fig. 3. The initial state of system is $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$.
In Fig. 5, we demonstrate the fidelity dynamics of the relevant states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ for various $\lambda$. It is found that while the state $|5\rangle_{a}|0\rangle$ is slightly populated, the population on the state $|0\rangle_{a}|5\rangle_{b}$ decreases monotonically with time, with a decay rate roughly proportional to $\lambda$ for $t\leq 0.3\tau$. In particular, when $t=0.1\tau$, we have $\mathcal{F}_{0,5}=0.713$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.601$ for $\lambda=1$ and $\mathcal{F}_{0,5}=0.557$ for $\lambda=1.2$. When $t=0.3\tau$, $\mathcal{F}_{0,5}=0.254$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.212$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.199$ for $\lambda=1.2$. When $t\geq 0.8\tau$, the population of the state $|0\rangle_{a}|5\rangle_{b}$ is almost vanishing for various $\lambda$. This result can be verified by the inset of Fig. 5, where the transition coefficients demonstrate a dramatic non-reciprocal relation $S_{21}(\tau)\gg S_{12}(\tau)$ when $\lambda>0.5$, justifying the unidirectional perfect absorbtion in the cavity-magnonic system. Note for $\lambda=0$, $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]=0$ or $S_{21}(\tau)=S_{12}(\tau)$ implies the disappearance of the non-reciprocal behavior or the bidirectional perfect state transfer in the Hermitian bosonic system Jin and Jing (2025e). It is consistent with the fact that under the condition of Eq. (35) with $\lambda=0$, the non-Hermitian Hamiltonian in Eq. (37) will become a Hermitian one Jin and Jing (2025e), i.e., $H(t)=iJ(t)a^{\dagger}b-iJ(t)b^{\dagger}a$.
@@ V Conclusion
In summary, we propose a versatile method for manipulating the general bosonic system through a non-Hermitian and time-dependent Hamiltonian. In sharp contrast to the conventional treatments of non-Hermitian systems, our theory is built upon the differential manifolds (instantaneous frames) rather than the system’s energy spectra, making the system dynamics independent of $\mathcal{PT}$-symmetry of Hamiltonian and the presence or absence of exceptional points.
In practice, the system dynamics can be described in the ancillary representations. By rotating from the time-dependent to the stationary ancillary representations, a purely geometric gauge potential emerges to shape the time evolution of the system, which is uniquely determined by the formation of ancillary operators. In the stationary representation, the Hamiltonian’s coefficient matrix can be upper triangularized under the triangularization condition with the projection operator, which suffices to activate two Heisenberg passages for flexible and perfect state transfers. Along the transitionless passages, arbitrary initial state can be perfectly transferred between the cavity and magnon modes without artificial normalization, including but not limited to Fock states, coherent states, the superposition of Fock states, the cat states, and even thermal states. The resulting passages do not show a clear relation with both $\mathcal{PT}$-symmetry of Hamiltonian and EPs in the energy spectrum. Moreover, the cavity-magnonic system can be used as a unidirectional perfect absorber under certain conditions. This work essentially extends our universal quantum control to cover the non-Hermitian continuous-variable systems.
@@ Acknowledgments
We acknowledge grant support from the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
@@ Appendix A A general recipe for ancillary operators
This appendix provides a brief recipe for constructing the ancillary basis operators $\{\mu_{k}(t)\}$ for a general bosonic systems of $N$ modes as well as the $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}$ in Eq. (3). Rooted in the geometric structure of the manifold of $\mu_{k}(t)$’s, the adjoint matrix $\mathcal{M}^{\dagger}(t)$ admits the representation Jin and Jing (2025b, e):
@@
With the definition in Eqs. (3) and (42), the rotation from the ancillary operators $\{\mu_{k}(t)\}$ to their stationary version $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$ can be performed by Jin and Jing (2025e)
@@
Appendix B Derivation of Eq. (21)
This appendix provides a detailed derivation of the non-Hermitian Hamiltonian (21) from the Lindblad master equation. In general, the dynamics of an open cavity magnonic system that is coupled to the traveling waves Metelmann and Clerk (2015); Wang et al. (2019a) can be described by
@@
where $H_{\rm coh}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+[J\exp(i\varphi)a^{\dagger}b+{\rm H.c.}]$ represents the eigen-energies of the two modes and the coherent coupling between them. The Lindblad superoperators are defined as $\mathcal{L}[o]\rho=o\rho o^{\dagger}-\{o^{\dagger}o,\rho\}/2$, $o=c,a,b$. The first superoperator is about $c\equiv ua+\exp{(i\Theta)}vb$ with coefficients $u$ and $v$. It implies the cooperative interactions between the two modes and the traveling wave with a damping rate $\eta$. $\Theta=0$ or $\Theta=\pi$ is determined by the propagation direction of the traveling waves. The second and the third superoperators are associated with the individual dissipations of the modes $a$ and $b$, with the damping rates $\beta$ and $\chi$, respectively.
@@
where the Hermitian conjugate superoperator $\mathcal{L}^{\dagger}[o]\equiv o^{\dagger}\rho o-\{o^{\dagger}o,\rho\}/2$, and the derivation from the second line to the third one has used the cyclic property of the density trace. According to Eq. (48), the time derivative of $\mathcal{O}_{H}(t)$ can be expressed as

### v2

@@ IV.2 Unidirectional perfect absorber
In this section, we show that the cavity-magnonic system under our UQC theory in Sec. IV.1 can simulate a nonreciprocal behavior under the same control conditions as in Fig. 3 for $\mu_{1}^{\dagger}(t)$ in the ket space. For example, when the Fock state is initially prepared in the magnon mode, i.e., $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$, the total population of the system is found to be nearly vanishing at the end of time evolution. Note the description about such a phenomenon is meaningful by our theory with no artificial renormalization. To a certain degree, this unidirectional absorbtion can be interpreted by the coupled-mode theory Xia et al. (2025) about coherent perfect absorber (CPA) Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025).
@@
$=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}\begin{pmatrix}-i\gamma_{a}&iJ(t)+i\Gamma\\ -iJ(t)+i\Gamma&-i\gamma_{a}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$.
@@
where $S_{21}(\tau)\approx 4\gamma_{1}J(\tau)/\gamma_{a}^{2}\neq 0$. Here the second equivalence holds when $\gamma_{1}=\gamma_{a}/2$. Equation (41) indicates that for a nonvacuum initial state in the magnon mode $b$, the total system population can become almost zero at a desired moment, i.e., $a(\tau),b(\tau)]^{T}=S(t=\tau)[0,b(0)]^{T}\approx(0,0)^{T}$. In practice, the non-Hermitian cavity magnonic system now becomes a promising candidate Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) for the unidirectional perfect absorber.
Figure 5: Fidelity dynamics about the individual states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ with various $\lambda$, the coupling strength or the dissipation rate of the system, under a $\mathcal{PT}$-symmetric-broken Hamiltonian. Inset: Numerical results for the logarithm of the nonreciprocity $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]$ as a function of $\lambda$. The other parameters are the same as Fig. 3. The initial state of system is $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$.
In Fig. 5, we demonstrate the fidelity dynamics of the relevant states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ for various $\lambda$. It is found that although the state $|5\rangle_{a}|0\rangle$ is slightly populated, the population on the state $|0\rangle_{a}|5\rangle_{b}$ decreases monotonically with time, with a decay rate roughly proportional to $\lambda$ for $t\leq 0.3\tau$. In particular, when $t=0.1\tau$, we have $\mathcal{F}_{0,5}=0.713$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.601$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.557$ for $\lambda=1.2$. When $t=0.3\tau$, $\mathcal{F}_{0,5}=0.254$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.212$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.199$ for $\lambda=1.2$. When $t\geq 0.8\tau$, the population of the state $|0\rangle_{a}|5\rangle_{b}$ is almost vanishing for various $\lambda$. This result can be verified by the inset of Fig. 5, where the transition coefficients demonstrate a dramatic non-reciprocal relation $S_{21}(\tau)\gg S_{12}(\tau)$ when $\lambda>0.5$, justifying the unidirectional perfect absorbtion in the cavity-magnonic system. Note for $\lambda=0$, $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]=0$, i.e., $S_{21}(\tau)=S_{12}(\tau)$ implies the disappearance of the non-reciprocal behavior, reminiscent of the fact that under the condition of Eq. (35) with $\lambda=0$, the non-Hermitian Hamiltonian in Eq. (37) becomes a Hermitian one Jin and Jing (2025g), i.e., $H(t)=iJ(t)a^{\dagger}b-iJ(t)b^{\dagger}a$. Our results then encompass the bidirectional perfect state transfer in the Hermitian bosonic system as a special case.
@@ V Conclusion
In summary, we propose a systematic and versatile method for manipulating the general bosonic system through a non-Hermitian and time-dependent Hamiltonian. In sharp contrast to conventional protocols, our theory describes the dynamics of non-Hermitian continuous-variable systems in the full Hilbert space through Heisenberg equations, does not ignore the quantum-jump terms, and relies on the gauge freedom of differential manifolds (instantaneous frames) rather than the eigen-spectrum. Consequently, beside realizing unidirectional perfect absorption, the control protocol enables perfect and flexible transfer of arbitrary states between bosonic systems, independent of $\mathcal{PT}$-symmetry of Hamiltonian or coefficient matrix and the presence or absence of exceptional points.
In practice, the system dynamics can be described in the ancillary representations. By rotating from the time-dependent to the stationary ancillary representations, a purely geometric gauge potential emerges to shape the time evolution of the system, which is uniquely determined by the formation of ancillary operators. In the stationary representation, the Hamiltonian’s coefficient matrix can be upper triangularized under Schur’s theorem with the projection operator, which suffices to activate two Heisenberg passages for flexible and perfect state transfers. Along the transitionless passages, arbitrary initial state can be perfectly transferred between the cavity and magnon modes without artificial normalization, including but not limited to Fock states, coherent states, the superposition of Fock states, the cat states, and even thermal states. The resulting passages do not show a clear relation with both $\mathcal{PT}$-symmetry and EPs. Moreover, the cavity-magnonic system can be used as a unidirectional perfect absorber under certain conditions. Essentially, this work extends the framework of our UQC theory, and provides a powerful tool for controlling the non-Hermitian continuous-variable systems.
@@ Acknowledgments
We acknowledge grant support from the National Natural Science Foundation of China (Grant No. U25A20199) and the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
@@ Appendix A A general recipe for ancillary operators
This appendix provides a brief recipe for constructing the ancillary basis operators $\{\mu_{k}(t)\}$ for a general bosonic systems of $N$ modes as well as the $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ in Eq. (3). Rooted in the geometric structure of the manifold of $\mu_{k}(t)$’s, the adjoint matrix $\mathcal{M}^{\dagger}(t)$ admits the representation Jin and Jing (2025c, g):
@@
With the definition in Eqs. (3) and (42), the rotation from the ancillary operators $\{\mu_{k}(t)\}$ to their stationary version $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$ can be performed by Jin and Jing (2025g)
@@
Appendix B Derivation of non-Hermitian Hamiltonian
This appendix provides a detailed derivation of the non-Hermitian Hamiltonian (20) from the Lindblad master equation. We emphasize that the quantum jump terms have been fully retained throughout the derivation Metelmann and Clerk (2015); Wang et al. (2019a) leading to Eq. (20), in contrast to those approaches only regarding the nonunitary evolution part after post-selection Han et al. (2024). In general, the dynamics of an open cavity magnonic system that is coupled to the traveling waves Metelmann and Clerk (2015); Wang et al. (2019a) can be described by
@@ Appendix B Derivation of non-Hermitian Hamiltonian
where $H_{\rm coh}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+[J\exp(i\varphi)a^{\dagger}b+{\rm H.c.}]$ represents the eigen-energies of the two modes and the coherent coupling between them. The Lindblad superoperators are defined as $\mathcal{L}[o]\rho=o\rho o^{\dagger}-\{o^{\dagger}o,\rho\}/2$, $o=c,a,b$. The first superoperator is about $c\equiv ua+\exp{(i\Theta)}vb$ with weight factors $u$ and $v$. It implies the cooperative interactions between the two modes and the traveling wave with a damping rate $\eta$. $\Theta=0$ or $\Theta=\pi$ is determined by the propagation direction of the traveling waves. The second and the third superoperators are associated with the individual dissipations of the modes $a$ and $b$, with the damping rates $\beta$ and $\chi$, respectively.
@@
where the Hermitian conjugate superoperator $\mathcal{L}^{\dagger}[o]\equiv o^{\dagger}\rho o-\{o^{\dagger}o,\rho\}/2$. The derivation from the second line to the third one in Eq. (48) has used the cyclic property of the density trace. Accordingly, the time derivative of $\mathcal{O}_{H}(t)$ can be expressed as